AN EMPTY PLASTIC BOX OF MASS m IS FOUND TO ACCELERATE UP AT THE RATE OF g/6 WHEN PLACED DEEP INSIDE WATER.HOW MUCH SAND SHOULD BE PUT INSIDE THE BOX SO THAT IT MAY ACCELERATE DOWN AT THE RATE OF g/6?

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

buoyant force=mg+mg/6=7mg/6,

let 'w' be mass of sand. (w+m)g-7mg/6=(w+m)g/6, solve for w